OpenStudy (anonymous):

Evaluate the function for the given value of x. f(x) = -6x2 - 4x - 4 , x = - 2 im stuck

OpenStudy (anonymous):

f(x) = -6x2 - 4x - 4=-6x^2-4x-4 f(-2)=-6(-2)^2-4(-2)-4 =-6(4)+8-4 =-24+4 =-20

OpenStudy (anonymous):

thanks metal for u

OpenStudy (anonymous):

welcome